Mathematics 241 – Multivariable Calculus Quiz 8 – November 15, 2013

Your Name:

A) (15) Let R be the triangular region in the xy-plane with corners at (0, 0), (1, 1) and (2, 0). Draw the region and set up the limits of integration for ∫ ∫ x dx dy to integrate x over R. Evaluate your iterated integral.

B) (5) Now set up an iterated integral (or integrals) of the form ∫ ∫ x dy dx to integrate over R. You do not need to evaluate this time.